# Polynomial-time isomorphism test for $k$-generated extensions of abelian groups

### v1

@@ Abstract
The group isomorphism problem asks whether two finite groups given by their Cayley tables are isomorphic or not. Although there are polynomial-time algorithms for some specific group classes, the best known algorithm for testing isomorphism of arbitrary groups of order $n$ has time complexity $n^{O(\log n)}$. We consider the group isomorphism problem for some extensions of abelian groups by $k$-generated groups for bounded $k$. In particular, we prove that one can decide isomorphism of abelian-by-cyclic extensions in polynomial time, generalizing a 2009 result of Le Gall for coprime extensions. As another application, we give a polynomial-time isomorphism test for abelian-by-simple group extensions, generalizing a 2017 result of Grochow and Qiao for central extensions. The main novelty of the proof is a polynomial-time algorithm for computing the unit group of a finite ring, which might be of independent interest.
@@ 1 Introduction
There have been many improvements for the group isomorphism problem for particular group classes, see, for example, the overview in [9]. Babai, Codenotti and Qiao showed [2] that isomorphism of groups with no normal abelian subgroups can be tested in polynomial time, in particular, their result highlights the importance of studying the group isomorphism problem for extensions of abelian groups. To fix our terminology, we say that a group $G$ with a normal subgroup $A$ is an extension of $A$ by $G/A$; we say that an extension is coprime, if $|A|$ is coprime to $|G/A|$. Le Gall proved [15] that the group isomorphism problem is solvable in polynomial time for coprime abelian-by-cyclic extensions, Qiao, Sarma and Tang [17] generalized that result to coprime extensions of abelian groups by $k$-generated groups for bounded $k$, where time complexity of the algorithm depended on $k$. The most general result of this flavor is by Babai and Qiao [4] which gives a polynomial isomorphism test for groups with abelian Sylow towers of normal subgroups; the key step in their algorithm resolves the isomorphism problem for certain coprime extensions of abelian groups [4, Theorem 1.2].
The main reason why there has been so much progress for coprime extensions is the fact that such extensions are split: if $G$ is a coprime extension of $A$ by $G/A$, then by Schur–Zassenhaus theorem there exists a subgroup $H\leq G$, $H\simeq G/A$, such that $G$ is a semidirect product $A\rtimes H$. Moreover, when $A$ is abelian, one can often utilize results from representation theory to study the action of $H$ on $A$. For example, by [4, Theorem 1.2], if $A$ is abelian and the full automorphism group of $H$ is known, then the full automorphism group of the coprime extension $A\rtimes H$ can be computed in polynomial time, which essentially reduces the problem for the original group $G$ to that problem for $H$.
@@
If we have two isomorphisms $\Phi_{1},\Phi_{2}:G\to G_{0}$ as in Theorem 1.1, then $\Phi_{1}\Phi_{2}^{-1}:G\to G$ is an automorphism of $G$ which stabilizes $A$ and acts trivially on $G/A$. In particular, the set of all isomorphisms $\Phi:G\to G_{0}$ inducing $\psi$ between $G/A$ and $G_{0}/A_{0}$ can be represented by a coset $\mathrm{Aut}_{0}(G,A)\cdot\Phi_{1}$, where $\mathrm{Aut}_{0}(G,A)$ is a subgroup of $\operatorname{Aut}(G)$ stabilizing $A$ and acting trivially on $G/A$.
@@
In Theorem 1.1 subgroups $A,A_{0}$ and the isomorphism $\psi:G/A\to G_{0}/A_{0}$ are given as a part of the input. In order to turn Theorem 1.1 into a proper isomorphism test, we need to be able to find suitable $A,A_{0}$ and $\psi$ in polynomial time. The following result covers an important special case when this is possible. The key property is that if $G/A$ is $k$-generated, then all isomorphisms from $G/A$ to $G_{0}/A_{0}$ can be listed in time polynomial in $n^{k}$.
@@
The main novelty which allowed us to prove Theorem 1.1 is the following result on the computation of unit groups of finite rings, which might be of independent interest.
@@
Our proof of Theorem 1.6 follows the proof of a result by Hofmann [11]. It was shown in [11, Theorem A] that the problem of computing the presentation of the unit group of a finite ring is polynomial-time equivalent to the factoring and discrete logarithm problems, for which no polynomial-time algorithm is known. It turns out that this difficulty can be overcome in the case when the finite ring in question is a subring of the full endomorphism ring of a finite abelian group given by its Cayley table.
@@ 2 Preliminaries
Let $G\leq\operatorname{Sym}(m)$ be a permutation group of degree $m$. We say that $G$ is given by permutation generators, if we are given a subset $S$ of $G$ which generates $G$. Note that any permutation from $\operatorname{Sym}(m)$ can be specified by a list of $m$ elements, so it occupies $O(m\log m)$ space. Hence, if $G$ is given to an algorithm by permutation generators, then the algorithm works in polynomial time if it works in time polynomial in $|S|\cdot m\cdot\log m$. It is well-known, see for instance [20, Exercise 4.1], that any generating set $S$ can be reduced to at most $m^{2}$ generators in polynomial time. Therefore we may safely assume that an algorithm on permutation groups works in polynomial time if it works in time polynomial in the degree of the permutation group. Subgroups of permutation groups are also specified by their generators. We refer the reader to [20, Section 3.1] for the library of polynomial-time algorithms for permutation groups, note, in particular, that tasks like finding orbits and pointwise stabilizers can be performed in polynomial time.
@@
Let $A$ be a finite abelian group. It is well-known that $A$ is isomorphic to a direct sum of $t$ cyclic groups of orders $n_{1},\dots,n_{t}$. We say that $A$ is given by its cyclic generators, if the algorithm is given a $t$-tuple $(n_{1},\dots,n_{t})$ of integers as input. The length of the input is $l=O(\sum_{i=1}^{t}\log n_{t})$, and elements of $A$ can be represented by $t$-tuples $(x_{1},\dots,x_{t})$, where $x_{i}\in\mathbb{Z}_{n_{i}}$, $i=1,\dots,t$. If $e_{i}=(0,\dots,1,\dots,0)$ is a generator of the $i$-th direct summand $\mathbb{Z}_{n_{i}}$, $i=1,\dots,t$, then clearly $e_{1},\dots,e_{t}$ generate $A$. With some abuse of terminology, we say that $e_{1},\dots,e_{t}$ are the cyclic generators of $A$.
@@
Subgroups of $A$ will be specified by their generating sets; we note that any set of generators of a subgroup can be reduced to a set of size polynomial in $l$ in polynomial time. Homomorphisms from $A$ into some other group are specified by listing images of $e_{1},\dots,e_{t}$ under the homomorphism in question. We refer the reader to [8, Chapter 2] for a list of polynomial-time algorithms for abelian groups given by cyclic generators. Observe also, that if the group $A$ is given by its Cayley table, then in time polynomial in $|A|$ one can compute a representation of $A$ by cyclic generators together with an explicit isomorphism from $A$ to $\mathbb{Z}_{n_{1}}\oplus\dots\oplus\mathbb{Z}_{n_{t}}$, see, for instance, [13].
@@
Let $R$ be a finite (not necessarily commutative) unitary ring, and let $R^{+}$ denote its additive group. We say that $R$ is given by additive generators, if the abelian group $R^{+}\simeq\mathbb{Z}_{n_{1}}\oplus\dots\oplus\mathbb{Z}_{n_{t}}$ is given by its cyclic generators and the ring multiplication $\cdot:R\times R\to R$ is specified by $t^{3}$ numbers $\alpha_{ijk}$ such that $e_{i}\cdot e_{j}=\sum_{k=1}^{t}\alpha_{ijk}e_{k}$, where $e_{1},\dots,e_{t}$ are the cyclic generators of $R^{+}$. The length of the input in this representation can be bounded by $O(t^{2}\cdot\sum_{i=1}^{t}\log n_{t})$. With some abuse of terminology, we will say that a subring or an ideal $I$ in $R$ is given by additive generators, if we are given a set of elements of $R$ which spans $I$ additively. Homomorphisms between rings are represented as homomorphisms between underlying abelian groups. We refer the reader to [8, Chapter 3] for some basic polynomial-time algorithms for finite rings. In particular, we note that tasks like computing a quotient of a ring by its ideal, and computing sums and products of ideals can be performed in polynomial time.
@@
Let $A\simeq\mathbb{Z}_{n_{1}}\oplus\dots\oplus\mathbb{Z}_{n_{t}}$ be an abelian group given by its cyclic generators $a_{1},\dots,a_{t}$, and let $R$ be a finite unitary ring given by its additive generators $e_{1},\dots,e_{m}$. If $A$ has a structure of a right $R$-module, then in order to give the module action of $R$ on $A$ to an algorithm, we specify $t^{2}\cdot m$ numbers $\beta_{ijk}$ such that $a_{i}^{e_{j}}=\sum_{k=1}^{t}\beta_{ijk}a_{k}$; here $a^{x}$ denotes the image of $a\in A$ under the action of $x\in R$. The length of the input in this case is bounded by $O(tm\cdot\sum_{i=1}^{t}\log n_{t})$. Homomorphisms between modules are represented as homomorphisms between underlying groups. The following result shows that isomorphism of finite modules can be determined in polynomial time.
@@ 3 Isomorphisms of extensions
Given a group word $w(t_{1},\dots,t_{k})$ in variables $t_{1},\dots,t_{k}$, we write $w(\overline{t})$ as a shorthand. Similarly, if $g_{1},\dots,g_{k}$ are elements of some group, we write $w(\overline{g})$ for $w(g_{1},\dots,g_{k})$.
@@ 4 Computing unit groups of finite rings
Let $A$ be a finite abelian group given by its Cayley table. Let $\operatorname{End}(A)$ be the ring of endomorphisms of $A$. We can encode every element $f\in\operatorname{End}(A)$ as a tuple $(f(a))_{a\in A}$ of $|A|$ elements of $A$. The size of each element in this representation is polynomial in $|A|$, and we can compute element sums and products (compositions) in polynomial time. Note that in general the order of $\operatorname{End}(A)$ is not polynomial in $|A|$, but we have a bound $|\operatorname{End}(A)|\leq|A|^{\log|A|}$.
Let $\operatorname{End}(A)^{+}$ denote the additive group of $\operatorname{End}(A)$, and let $e_{1},\dots,e_{t}$ be cyclic generators of $A$. For each $a\in A$ set $N(a)=\{b\in A\mid b^{|a|}=1\}$. For $i=1,\dots,t$ and $a\in N(e_{i})$ we can define an endomorphism $f_{i,a}$ of $A$ by $f_{i,a}(x_{1}e_{1}+\dots+x_{t}e_{t})=x_{i}a$, where $x_{1},\dots x_{t}\in\mathbb{Z}$. Note that $f_{i,a}$, $i=1,\dots,t$, $a\in N(e_{i})$, generate $\operatorname{End}(A)^{+}$ additively. The exponent of $\operatorname{End}(A)^{+}$ is at most $|A|$, and hence one can compute the cyclic generators of $\operatorname{End}(A)^{+}$ and express any given element in terms of these generators in time polynomial in $|A|$.
Let $R$ be a unitary subring of $\operatorname{End}(A)$ given by its additive generators. Let $R^{\times}$ denote the unit group of $R$, and note that $R^{\times}$ acts on $A$ by automorphisms, in particular, every element of $R^{\times}$ is a permutation on $A$. Hence we can specify $R^{\times}$ as a permutation group in terms of its permutation generators; their number is polynomial in $|A|$.
Recall that if $J$ is a nilpotent ideal of $R$, then every element of $1+J=\{1+x\mid x\in I\}$ is invertible, so $1+J$ is a subgroup of $R^{\times}$ and can also be specified by (polynomially many) permutation generators.
@@
Let $A$ be a finite abelian group given by its Cayley table. If $R$ is a unitary subring of $\operatorname{End}(A)$ given by additive generators $S$, and $J$ is a nilpotent ideal of $R$ given by additive generators $T$, then we can compute permutation generators of $1+J$ in time polynomial in $|A|$, $|S|$ and $|T|$.
@@
Note that we can compute additive generators of $I^{2}$ in polynomial time by considering all pairwise products of the additive generators of $I$. Now, observe that the map $\phi:1+I\to I/I^{2}$ given by $\phi(1+x)=x+I^{2}$, $x\in I$, is a homomorphism from the multiplicative group $1+I$ to the additive group $I/I^{2}$. The kernel of $\phi$ is $1+I^{2}$, and note that for any $g\in I/I^{2}$ we can compute some preimage in $1+I$ in polynomial time. We can compute at most $\log|I/I^{2}|\leq\log|R|$ additive generators of $I/I^{2}$ in polynomial time, and let $1+g_{i}$, $i=1,\dots,r$, be preimages in $1+I$ of these generators. Since $|R|\leq|A|^{\log|A|}$ and thus $\log|R|\leq(\log|A|)^{2}$, our algorithm works in time polynomial in $|A|$ and the number of generators of $R$ and $I$.
Returning to the original task, let $n$ be the minimal positive integer such that $J^{n}=0$. Set $k=\lceil\log(n)\rceil$ and $J_{i}=J^{2^{k-i}}$, $i=0,\dots,k$. We have a series of normal subgroups $N_{i}=1+J_{i}$, $i=0,\dots,k$, where $N_{0}=1$ and $N_{k}=1+J$. By the argument from the previous paragraph applied to each quotient $N_{i+1}/N_{i}$, $i=0,\dots,k-1$, we obtain generators of the group $N_{k}$. Since $k\leq 1+\log|R|\leq 1+(\log|A|)^{2}$, the algorithm works in polynomial time. ∎
@@
Let $\mathbb{F}$ be a finite field of characteristic $p$ and order $p^{f}$, $f\geq 1$. Since $\mathbb{F}$ is a ring, we can specify it to our algorithms by additive generators. In particular, field operations in $\mathbb{F}$ can be performed in time polynomial in $f$ and $\log p$. A $n\times n$ matrix over $\mathbb{F}$ is encoded directly in the algorithm (as a matrix of elements from $\mathbb{F}$) and hence has size polynomial in $n,f$ and $\log p$. We say that an isomorphism between some $M_{i}$ and a matrix ring $M_{n}(\mathbb{F})$ can be found explicitly, if given an element of $M_{i}$ we can compute the corresponding $n\times n$ matrix in polynomial time, and vice versa.
@@
Minimal ideals can be computed in polynomial time by [18, Corollary 3.2]. A polynomial-time algorithm for finding an explicit isomorphism from $M_{i}$ to $M_{n_{i}}(\mathbb{F}_{q_{i}})$ follows from [18, Theorem 5.2]. The algorithm in [18, Theorem 5.2] is an $f$-algorithm, i.e. it is a polynomial-time algorithm which has access to an oracle for factoring polynomials over finite fields. Since our algorithms are allowed to run in time polynomial in $d$ and $p$ (versus $\log p$), we can use Berlekamp’s factoring algorithm [5] to eliminate the use of the oracle. ∎
Now we are ready to prove the main result of this section and show that the unit group of a finite ring can be computed in polynomial time if the ring acts on a finite abelian group given by its Cayley table.
Proof of Theorem 1.6. Let $A$ be a finite abelian group given by its Cayley table, and let $R$ be a unitary subring of $\operatorname{End}(A)$ given by its additive generators. Note that all prime divisors of $|R|$ are bounded by $|A|$, so we can factor the order of $R$ into prime power factors in polynomial time. Given this factorization, decompose $R^{+}$ into a direct sum of abelian groups
$R^{+}=R_{1}\oplus\dots\oplus R_{k},$
where each $R_{i}$ has prime power order, $k\leq\log|R|\leq(\log|A|)^{2}$, and $|R_{i}|$ and $|R_{j}|$ are coprime for $i\neq j$. Observe that $R_{i}$, $i=1,\dots,k$, is a subring of $R$, and note that
$R^{\times}=R_{1}^{\times}\times\dots\times R_{k}^{\times}.$
@@
Without loss of generality we may assume that $R$ has order equal to a power of a prime $p$. Let $J$ be the Jacobson radical of $R$. There is an exact sequence of groups
@@
where each $M_{i}$, $i=1,\dots,k$, is isomorphic to a full matrix ring $M_{n_{i}}(q_{i})$. This decomposition and explicit isomorphisms between $M_{i}$ and $M_{n_{i}}(q_{i})$ can be found in polynomial time by Lemma 4.2. Since
@@
and $M_{i}^{\times}\simeq\operatorname{GL}_{n_{i}}(q_{i})$, $i=1,\dots,k$, it suffices to find multiplicative generators of $\operatorname{GL}_{n_{i}}(q_{i})\leq M_{n_{i}}(q_{i})$. Indeed, once we compute the multiplicative generators of $M_{i}^{\times}$, $i=1,\dots,k$, we can lift them back to $R$ and obtain the generators of $R^{\times}$ as wanted.
@@
it suffices to find generators of the multiplicative group of a finite field $\mathbb{F}_{q_{i}}$. By [16, Section 2], in polynomial time we can find an irreducible polynomial $f\in\mathbb{Z}_{p}[x]$ and an explicit isomorphism between $\mathbb{F}_{q_{i}}$ and the quotient-ring $\mathbb{Z}_{p}[x]/(f)$. By [21, Theorem 1.1], in time polynomial in $\deg f=\log q_{i}/\log p$ and $p$ we can find a subset of elements of $\mathbb{Z}_{p}[x]/(f)$ (of polynomial size) which contains a primitive root, in particular, we can compute multiplicative generators of $\mathbb{F}_{q_{i}}^{\times}$, and hence of $D_{n_{i}}(q_{i})$, in polynomial time.

### v2

@@ Abstract
The group isomorphism problem asks whether two finite groups given by their Cayley tables are isomorphic or not. Although there are polynomial-time algorithms for some specific group classes, the best known algorithm for testing isomorphism of arbitrary groups of order $n$ has time complexity $n^{O(\log n)}$. We consider the group isomorphism problem for some extensions of abelian groups by $k$-generated groups for bounded $k$. In particular, we prove that one can test isomorphism of abelian-by-cyclic extensions in polynomial time, generalizing a 2009 result of Le Gall for coprime extensions. As another application, we give a polynomial-time isomorphism test for abelian-by-simple group extensions, generalizing a 2017 result of Grochow and Qiao for central extensions. The main novelty of the proof is a polynomial-time algorithm for computing the unit group of a finite ring, which might be of independent interest.
@@ 1 Introduction
There have been many improvements for the group isomorphism problem for particular group classes, see, for example, the overview in [9]. Babai, Codenotti and Qiao showed [2] that isomorphism of groups with no normal abelian subgroups can be tested in polynomial time, in particular, their result highlights the importance of studying the group isomorphism problem for extensions of abelian groups. To fix our terminology, we say that a group $G$ with a normal subgroup $A$ is an extension of $A$ by $G/A$; we say that an extension is coprime, if $|A|$ is coprime to $|G/A|$. Le Gall proved [15] that the group isomorphism problem is solvable in polynomial time for coprime abelian-by-cyclic extensions, Qiao, Sarma and Tang [17] generalized that result to coprime extensions of abelian groups by $k$-generated groups for bounded $k$, where time complexity of the algorithm depended on $k$ exponentially. The most general result of this flavor is by Babai and Qiao [4] which gives a polynomial isomorphism test for groups with abelian Sylow towers of normal subgroups; the key step in their algorithm resolves the isomorphism problem for certain coprime extensions of abelian groups [4, Theorem 1.2].
One of the reasons why there has been so much progress for coprime extensions is the fact that such extensions are split: if $G$ is a coprime extension of $A$ by $G/A$, then by Schur–Zassenhaus theorem there exists a subgroup $H\leq G$, $H\simeq G/A$, such that $G$ is a semidirect product $A\rtimes H$. Moreover, when $A$ is abelian, one can often utilize results from representation theory to study the action of $H$ on $A$. For example, by [4, Theorem 1.2], if $A$ is abelian and the full automorphism group of $H$ is known, then the full automorphism group of the coprime extension $A\rtimes H$ can be computed in polynomial time, which essentially reduces the problem for the original group $G$ to that problem for $H$.
@@
If we have two isomorphisms $\Phi_{1},\Phi_{2}:G\to G_{0}$ as in Theorem 1.1, then $\Phi_{1}\Phi_{2}^{-1}:G\to G$ is an automorphism of $G$ which stabilizes $A$ and acts trivially on $G/A$. In particular, the set of all isomorphisms $\Phi:G\to G_{0}$ inducing $\psi$ between $G/A$ and $G_{0}/A_{0}$ can be represented by a coset $\mathrm{Aut}_{0}(G,A)\cdot\Phi_{1}$, where $\mathrm{Aut}_{0}(G,A)$ is a subgroup of $\operatorname{Aut}(G)$ consisting of all elements which stabilize $A$ and act trivially on $G/A$.
@@
In Theorem 1.1 the subgroups $A,A_{0}$ and isomorphism $\psi:G/A\to G_{0}/A_{0}$ are given as a part of the input. In order to turn Theorem 1.1 into a proper isomorphism test, we need to be able to find suitable $A,A_{0}$ and $\psi$ in polynomial time. The following result covers an important special case when this is possible. The key property is that if $G/A$ is $k$-generated, then all isomorphisms from $G/A$ to $G_{0}/A_{0}$ can be listed in time polynomial in $n^{k}$.
@@
The main novelty which allowed us to prove Theorem 1.1 is the following result on the computation of unit groups of finite rings, which might be of independent interest. The way rings are encoded in our algorithms is explained in detail in Section 2.
@@
Let $R$ be a finite unitary ring given by its additive generators. If $p_{\max}$ is the largest prime divisor of $|R|$, then we can compute the multiplicative generators of $R^{\times}$ in time polynomial in $\log|R|$ and $p_{\max}$.
Our proof of Theorem 1.6 follows the proof of a result by Hofmann [11]. It was shown in [11, Theorem A] that the problem of computing an effective presentation of the unit group of a finite ring is probabilistic polynomial-time equivalent to the factoring and discrete logarithm problems, for which no polynomial-time algorithm is known. It turns out that this difficulty can be overcome if the time complexity of the algorithm is allowed to depend on the size of the largest prime divisor of the order of the ring. In our application to the group isomorphism problem $R$ will be a subring of the endomorphism ring of a finite abelian group $A$ given by its Cayley table, see Corollary 4.3. In particular, all prime divisors of $|R|$ will be bounded by $|A|$.
@@ 2 Preliminaries
Let $G\leq\operatorname{Sym}(m)$ be a permutation group of degree $m$. We say that $G$ is given by permutation generators, if we are given a subset $S$ of $G$ which generates $G$. Note that any permutation from $\operatorname{Sym}(m)$ can be specified by a list of $m$ elements, so it occupies $O(m\log m)$ space. Hence, if $G$ is given to an algorithm by permutation generators, then the algorithm works in polynomial time if it works in time polynomial in $|S|\cdot m\cdot\log m$. It is well-known, see for instance [20, Exercise 4.1], that any generating set $S$ can be reduced to at most $m^{2}$ generators in time polynomial in $m$ and $|S|$. Therefore we may safely assume that an algorithm on permutation groups works in polynomial time if it works in time polynomial in the degree of the permutation group. Subgroups of permutation groups are also specified by their generators. We refer the reader to [20, Section 3.1] for the library of polynomial-time algorithms for permutation groups, note, in particular, that tasks like finding orbits and pointwise stabilizers can be performed in polynomial time.
@@
Let $A$ be a finite abelian group. It is well-known that $A$ is isomorphic to a direct sum of $t$ cyclic groups of orders $n_{1},\dots,n_{t}$. We say that $A$ is given by its cyclic generators, if the algorithm is given a $t$-tuple $(n_{1},\dots,n_{t})$ of integers as input. The length of the input is $l=O(\sum_{i=1}^{t}\log n_{i})$, and elements of $A$ can be represented by $t$-tuples $(x_{1},\dots,x_{t})$, where $x_{i}\in\mathbb{Z}_{n_{i}}$, $i=1,\dots,t$. If $e_{i}=(0,\dots,1,\dots,0)$ is a generator of the $i$-th direct summand $\mathbb{Z}_{n_{i}}$, $i=1,\dots,t$, then clearly $e_{1},\dots,e_{t}$ generate $A$. With some abuse of terminology, we say that $e_{1},\dots,e_{t}$ are the cyclic generators of $A$. Observe also that $l=O(\log|A|)$.
@@
Subgroups of $A$ will be specified by their generating sets; we note that any set of generators $S\subseteq A$ of a subgroup can be reduced to a set of size at most $t$ in time polynomial in $|S|$ and $l$. Homomorphisms from $A$ into some other group are specified by listing images of $e_{1},\dots,e_{t}$ under the homomorphism in question. We refer the reader to [8, Chapter 2] for a list of polynomial-time algorithms for abelian groups given by cyclic generators. Observe also, that if the group $A$ is given by its Cayley table, then in time polynomial in $|A|$ one can compute a representation of $A$ by cyclic generators together with an explicit isomorphism from $A$ to $\mathbb{Z}_{n_{1}}\oplus\dots\oplus\mathbb{Z}_{n_{t}}$, see, for instance, [13].
@@
Let $R$ be a finite (not necessarily commutative) unitary ring, and let $R^{+}$ denote its additive group. We say that $R$ is given by additive generators, if the abelian group $R^{+}\simeq\mathbb{Z}_{n_{1}}\oplus\dots\oplus\mathbb{Z}_{n_{t}}$ is given by its cyclic generators and the ring multiplication $\cdot:R\times R\to R$ is specified by $t^{3}$ numbers $\alpha_{ijk}$ such that $e_{i}\cdot e_{j}=\sum_{k=1}^{t}\alpha_{ijk}e_{k}$, where $e_{1},\dots,e_{t}$ are the cyclic generators of $R^{+}$. The length of the input in this representation can be bounded by $l=O(t^{2}\cdot\sum_{i=1}^{t}\log n_{i})$. Note that $\log|R|$ is bounded polynomially in terms of $l$ and vice versa, $l=O((\log|R|)^{3})$. In particular, an algorithm working with a finite ring $R$ works in polynomial time if and only if the algorithm works in time polynomial in $\log|R|$.
With some abuse of terminology, we will say that a subring or an ideal $I$ in $R$ is given by additive generators, if we are given a set of elements of $R$ which spans $I$ additively. Homomorphisms between rings are represented as homomorphisms between underlying abelian groups. We refer the reader to [8, Chapter 3] for some basic polynomial-time algorithms for finite rings. In particular, we note that tasks like computing a quotient of a ring by its ideal, and computing sums and products of ideals can be performed in polynomial time.
@@
Let $A\simeq\mathbb{Z}_{n_{1}}\oplus\dots\oplus\mathbb{Z}_{n_{t}}$ be an abelian group given by its cyclic generators $a_{1},\dots,a_{t}$, and let $R$ be a finite unitary ring given by its additive generators $e_{1},\dots,e_{m}$. If $A$ has a structure of a right $R$-module, then in order to give the module action of $R$ on $A$ to an algorithm, we specify $t^{2}\cdot m$ numbers $\beta_{ijk}$ such that $a_{i}^{e_{j}}=\sum_{k=1}^{t}\beta_{ijk}a_{k}$; here $a^{x}$ denotes the image of $a\in A$ under the action of $x\in R$. The length of the input in this case is bounded by $O(tm\cdot\sum_{i=1}^{t}\log n_{i})$. Homomorphisms between modules are represented as homomorphisms between underlying groups. The following result shows that isomorphism of finite modules can be determined in polynomial time.
@@ 3 Isomorphisms of extensions
Given a group word $w(t_{1},\dots,t_{k})$ in variables $t_{1},\dots,t_{k}$, we write $w(\overline{t})$ as a shorthand. Similarly, if $g_{1},\dots,g_{k}$ are elements of some group $G$, we write $w(\overline{g})$ for $w(g_{1},\dots,g_{k})\in G$.
@@ 4 Computing unit groups of finite rings
Let $R$ be a finite unitary ring given by its additive generators, and let $R^{\times}$ denote its unit group. Our goal is to compute (polynomially many) elements of $R$, which generate $R^{\times}$ multiplicatively.
Recall that if $J$ is a nilpotent ideal of $R$, then every element of $1+J=\{1+x\mid x\in I\}$ is invertible, so $1+J$ is a subgroup of $R^{\times}$.
@@
If $R$ is a finite unitary ring given by additive generators and $J$ is a nilpotent ideal of $R$ given by additive generators $T$, then we can compute multiplicative generators of $1+J$ in time polynomial in $\log|R|$ and $|T|$.
@@
Note that we can compute additive generators of $I^{2}$ in polynomial time by considering all pairwise products of the additive generators of $I$. Now, observe that the map $\phi:1+I\to I/I^{2}$ given by $\phi(1+x)=x+I^{2}$, $x\in I$, is a homomorphism from the multiplicative group $1+I$ to the additive group $I/I^{2}$. The kernel of $\phi$ is $1+I^{2}$, and note that for any $g\in I/I^{2}$ we can compute some preimage in $1+I$ in polynomial time. We can compute at most $\log|I/I^{2}|\leq\log|R|$ additive generators of $I/I^{2}$ in polynomial time, and let $1+g_{i}$, $i=1,\dots,r$, be preimages in $1+I$ of these generators.
Returning to the original task, let $n$ be the minimal positive integer such that $J^{n}=0$. Set $k=\lceil\log(n)\rceil$ and $J_{i}=J^{2^{k-i}}$, $i=0,\dots,k$. We have a series of normal subgroups $N_{i}=1+J_{i}$, $i=0,\dots,k$, where $N_{0}=1$ and $N_{k}=1+J$. By the argument from the previous paragraph applied to each quotient $N_{i+1}/N_{i}$, $i=0,\dots,k-1$, we obtain generators of the group $N_{k}$; note that the total number of generators obtained is at most $k\cdot\log|R|$. Since $k\leq 1+\log|R|$, the algorithm works in polynomial time. ∎
@@
Let $\mathbb{F}$ be a finite field of characteristic $p$ and order $p^{f}$, $f\geq 1$. Since $\mathbb{F}$ is a ring, we can specify it to our algorithms by additive generators. In particular, field operations in $\mathbb{F}$ can be performed in time polynomial in $f$ and $\log p$. An $n\times n$ matrix over $\mathbb{F}$ is encoded directly in the algorithm (as a matrix of elements from $\mathbb{F}$) and hence has size polynomial in $n,f$ and $\log p$. We say that an isomorphism between some $M_{i}$ and a matrix ring $M_{n}(\mathbb{F})$ can be found explicitly, if given an element of $M_{i}$ we can compute the corresponding $n\times n$ matrix in polynomial time, and vice versa.
@@
Minimal ideals can be computed in polynomial time by [18, Corollary 3.2]. A polynomial-time algorithm for finding an explicit isomorphism from $M_{i}$ to $M_{n_{i}}(\mathbb{F}_{q_{i}})$ follows from [18, Theorem 5.2], see also [18, Corollary 4.7]. The algorithm in [18, Theorem 5.2] is an $f$-algorithm, i.e. it is a polynomial-time algorithm which has access to an oracle for factoring polynomials over finite fields. Since our algorithms are allowed to run in time polynomial in $d$ and $p$ (versus $\log p$), we can use Berlekamp’s factoring algorithm [5] to eliminate the use of the oracle. ∎
Now we are ready to prove the main result of this section and show that the unit group of a finite ring can be computed in polynomial time if the largest prime divisor of the order of the ring is considered part of the input.
Proof of Theorem 1.6. Let $R$ be a finite unitary ring given by its additive generators, and let $p_{\max}$ be the largest prime divisor of $|R|$. In time polynomial in $p_{\max}$ and $\log|R|$ we can compute $|R|$ and decompose it into a product of prime power factors. Given this factorization, decompose $R^{+}$ into a direct sum of abelian groups
$R^{+}=R_{1}\oplus\dots\oplus R_{s},$
where each $R_{i}$ has prime power order, $s\leq\log|R|$, and $|R_{i}|$ and $|R_{j}|$ are coprime for $i\neq j$. Observe that $R_{i}$, $i=1,\dots,s$, is a subring of $R$, and note that
$R^{\times}=R_{1}^{\times}\times\dots\times R_{s}^{\times}.$
@@
Without loss of generality we may assume that $|R|$ is a power of a prime $p$. Let $J$ be the Jacobson radical of $R$. There is an exact sequence of groups
@@
where each $M_{i}$, $i=1,\dots,k$, is isomorphic to a full matrix ring $M_{n_{i}}(q_{i})$, and $k\leq\log|R|$. This decomposition and explicit isomorphisms between $M_{i}$ and $M_{n_{i}}(q_{i})$ can be found in polynomial time by Lemma 4.2. Since
@@
and $M_{i}^{\times}\simeq\operatorname{GL}_{n_{i}}(q_{i})$, $i=1,\dots,k$, it suffices to find multiplicative generators of $\operatorname{GL}_{n_{i}}(q_{i})\leq M_{n_{i}}(q_{i})$. Indeed, once we compute the multiplicative generators of $M_{i}^{\times}$, $i=1,\dots,k$, we can lift them back to $R$ and together with generators of $1+J$ we obtain the generators of $R^{\times}$, as required.
@@
it suffices to find generators of the multiplicative group of a finite field $\mathbb{F}_{q_{i}}$. By [16, Section 2], in polynomial time we can find an irreducible polynomial $f\in\mathbb{Z}_{p}[x]$ and an explicit isomorphism between $\mathbb{F}_{q_{i}}$ and the quotient-ring $\mathbb{Z}_{p}[x]/(f)$. By [21, Theorem 1.1], in time polynomial in $\deg f=\log q_{i}/\log p$ and $p$ we can find a subset of elements of $\mathbb{Z}_{p}[x]/(f)$ (of polynomial size) which contains a primitive root, in particular, we can compute multiplicative generators of $\mathbb{F}_{q_{i}}^{\times}$, and hence of $D_{n_{i}}(q_{i})$, in polynomial time. Note that $\deg f\leq\log|R|$ and $p\leq p_{\max}$.